OpenStudy (anonymous):

Calculate the discriminant and use it to determine how many real-number roots the equation has. 3x^2 - 6x + 1 = 0 three real-number roots two real-number roots one real-number root no real-number roots

OpenStudy (anonymous):

what do you think it is?

OpenStudy (anonymous):

Lol I Dont Know :S

OpenStudy (anonymous):

can you give me the value of the discriminant?

OpenStudy (anonymous):

discriminant = b^2 - 4(a)(c).... in \(\large ax^2+bx+c=0 \)

OpenStudy (anonymous):

can you at least identify a= b= c= ???

OpenStudy (anonymous):

compare your equation with \(\large ax^2+bx+c=0 \) and match up the coefficients...

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

Oh Sorry This Is Not My Math Its My Sisters And Im Helping Her So I dont Know Much :/

OpenStudy (anonymous):

i think she'd learn better if you let her do her own work, don't you think? :)

OpenStudy (anonymous):

 since the number is positive, there are 2 real solutions.